# Gauging or extending bulk and boundary conformal field theories: Application to bulk and domain wall problem in topological matter and their descriptions by (mock) modular covariant

### v2

@@ Abstract
We formulate the $Z_{N}$ group gauging operations (or group extensions) in (smeared) boundary conformal field theories (BCFTs) and bulk conformal field theories and study their applications to various phenomena in topologically ordered systems. We apply the resultant theories to analyze the correspondence between the renormalization group (RG) flow of CFTs and the classification of topological quantum field theories with the testable information of general classes of partition functions. One can obtain the bulk topological properties of $2+1$ dimensional topological ordered phase corresponding to the massive RG flow of $1+1$ dimensional systems equivalent to the smeared BCFT. We present an obstruction of mass condensation for the smeared BCFT analogous to the Lieb-Shultz-Mattis theorem for noninvertible symmetry. Corresponding to the bulk topological degeneracies in $2+1$ dimensions and gapped quantum phases in $1+1$ dimensions, we construct a new series of BCFT.
We also study the implications of the massless RG flow of $1+1$ dimensional CFTs corresponding to the domain wall or coupling problem in $2+1$ dimensional topological orders which corresponds to the earlier proposal by L. Kong and H. Zheng in [Nucl. Phys. B 966 (2021), 115384], arXiv:1912.01760 closely related to the integer-spin simple current by Schellekens and Gato-Rivera. We reveal that two distinct kinds of massless RG flow exist, coming from anomaly cancellation and anomaly matching related to Witt equivalence of bulk CFTs. The anomaly cancellation flows produce a new series of modular invariants as coupled models and the anomaly matching flows produce nontrivial domain walls changing the chirality of anyons. Our paper gives a unified direction for the future theoretical and numerical studies of the topological phase based on the established data of classifications of conformal field theories or modular invariants (and the corresponding category theories).
@@
Anyon condensation and Witt equivalence are the most fundamental structures classifying topological orders (TOs) or corresponding topological quantum field theories (TQFTs) in contemporary physics and related mathematics[1, 2, 3]. The theories are deformed to each other without obstruction when the two theories are properly related by the classifications. Whereas the systematic studies on such classification schemes are still in development and sometimes contain controversies, the fundamental significance of abstract algebra of anyons and the corresponding category theories have already been established to some extent (see [4, 5, 6, 7] and reference therein, for example). Combining these ideas with gauging (or group extension) and orbifolding which has been studied in the context of renormalization group (RG) flow[8, 9, 10], one can expect some general correspondence between the classification of TOs in $2+1$ dimensions and that of the RG flows in $1+1$ dimensional CFTs[11, 12, 13, 14, 15]. This is a modern version of bulk-edge correspondence or correspondence between conformal field theory (CFT) and TQFT known as CFT/TQFT[16, 17].
@@
The symmetry algebraic or category-theoretic version of CFT/TQFT[18, 19, 20, 21, 22] [^1] We thank Liang Kong for notifying us a lot references and their status in the fields., called “topological holography" [24], has been studied widely recently (However, the terminology “topological holography" appeared earier in a different contenxt in [25] representing systems satisyfying both topological and holographic property). The resultant sandwich construction[26, 27, 28] has applied for several models[29, 30, 31, 32, 33, 34, 35, 36]. Based on these backgrounds, constructing a classification of CFTs is fundamentally important for the classification of TOs. However, we stress that in such generalized symmetry-based studies, the theoretical and mathematical framework on the famous TOs, fractional quantum Hall effects (FQHEs), have been still under development regardless of their experimental significance[37]. It might be surprising for readers, but the existing famous framework, modular-tensor-category (MTC) or bosonic chiral CFT (CCFT), does not correspond to the FQHEs in general. It is necessary to extend the bosonic theory by a finite group and this is called simple current extension[38, 39, 40, 41, 42, 43]. When concentrating our attention on the $Z_{2}$ or fermionic case, such extended structure has already appeared in the established works [17, 44, 45, 46, 47, 48]. More generally, one can see the corresponding arguments widely in older theoretical literature [45, 49, 50, 51, 52, 53, 54]. The resultant structure has captured attention of the fields as realizations of supersymmetry in condensed matter[55, 56, 57, 58, 59, 60, 61, 62, 63]. More recently, the nontriviality of the corresponding categorical frameworks has been studied in [64, 65, 66]. The main difficulty in studying group extensions of CFTs is the treatment of (fermionic) zero modes [67] (or semion[68]). This zero mode counting results in the degeneracies of wavefunctions of FQHEs[17, 69], but the semion structure has not been studied extensively except for several earlier works[67, 46, 44, 70, 71, 72] and the recent works by the author [61, 73, 12, 74, 32]. This is in parallel to the recent development of the studies on fermionic or gauged models[75, 76, 77, 78, 79, 80, 81, 82] (The references are huge, so we have only listed established literature or literature containing discussions on indicators). The zero modes are closely related to the order-disorder duality of lattice models[83], and the group extension is a procedure to unify order and disorder operators canonically in QFTs. Hence by including nonlocal objects, the theory after the extension becomes outside of MTCs. The resultant theory has been studied as a super-fusion category or premodular fusion category, for example, but this research direction is still in development.
@@
Figure 1: Summary of the relationship between RG flow in CFTs and the notions in TQFTs. The massive RG flow of CFTs or the (extended) smeared BCFTs[84] in this work determine the structure of bulk topological order with a close connection to the entanglement entropy[85, 86, 87, 88, 89]. In TQFT, the massive RG flow in CFTs corresponds to the gluing operation. On the other hand, the massless RG flow of CFTs implies the (Witt) equivalence relation between group extended TQFTs (or extended SFCs[90, 91]). Applying the interface or domain wall known as gapped or symmetry preserving domain wall[11] in TQFT corresponds to the massless RG flow. By analyzing $BCFT_{D}$ with $D\geq 3$ as ancillary CFT [92], a similar phenomenology will be true in general. We also note that the existing Witt equivalence for MTC (not SFC) seems too strong when applying to realistic settings[93, 6, 94, 95] (we stress that this does not mean their arguments are wrong. This implies the significance of extension for application to more realistic settings). One needs to introduce (or relax) the equivalence relation to the gauged bulk or chiral CFT.
@@
When studying classifications of CFTs, there exist two fundamental aspects, symmetry and renormalization group (RG) flow. Recently, it has been gradually established that a RG flow to a gapped phase can be mapped to the classification of boundary conformal field theory (BCFT)[96, 97, 98, 84], and a massless RG flow [99, 100, 101] can be mapped to the classification of the defect (or interface) of CFTs[102, 103] (We also note the literature on the massless RG flows [104, 105, 106, 107, 108, 109, 110, 111, 112]). In the context of CFT/TQFT, the former corresponds to the classification of bulk states or entanglement spectrum of TOs [85, 86, 87, 88, 89] and the latter corresponds to the classification of boundary anyons induced by the Witt equivalence [11, 12, 13, 14] (Fig. 1). Corresponding discussions for TQFT side can be seen in the studies of the gapped domain wall in [113, 114, 93, 19, 6, 20, 115, 21, 116] which can be interpreted as a particular class of symmetry preserving domain walls in [11, 12].
@@
The idea of RG domain wall is surprisingly concise because this is just attaching two systems connected under a massless RG flow [99, 100, 101] through a particular type of domain wall[102, 103]. Hence, one can map the problem of RG to an algebraic problem of domain wall (or conformal interface) between ultraviolet (UV) and infrared (IR) CFTs. Usually, the verification of RG requires extensive analytical calculations of quantum field theories (QFTs) and numerical tests in lattice models, and they tend to be difficult to formulate rigorously or mathematically. Moreover, by using a folding trick [117] to the system divided by an RG domain wall, one can map the problem to the product of RG-connected CFT with a boundary[102] and this is in the scope of BCFT analysis. Hence one can apply techniques in defect or boundary CFTs rigorously and systematically in principle (Fig.2). Moreover, the validity of the RG domain wall in the minimal CFTs has been tested more rigorously in [118] and one can expect the RG domain wall can provide a rigorous formulation of RGs. In this sense, further systematic studies on CFT with defects and boundaries are fundamentally important for the systematic classifications of TOs. We note recent works for the RG domain wall and related junction problems[119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 15, 133].
@@
Figure 2: Folding trick applied to coupled models. By applying the folding trick[117], one can obtain the correspondence between a coupled model and anyon transformation law of two theories. Similar arguments can be seen in [11, 116], for example. Because of the folding, the orientation or chirality of one of the two bulk CFTs is reversed. The coupled models in this work are the multicomponent or multilayered TOs. The mapping $D_{\text{RG}}$ corresponding to the symmetry preserving domain wall[11] or RG domain wall[103] classifies the TOs, possibly including FQHEs with nonchiral anyon. By shrinking one of the two systems, $CFT^{\prime}_{D}$ in the righthand side of the figure, one can obtain the mismatch of bulk theory $CFT_{D}$ and boundary theory $CFT^{\prime}_{D}$. This explains recent controversies on the thermal Hall conductance in experimental settings[134, 135].
@@
In this work, we systematically study algebraic properties of $2+1$ dimensional TOs as RG flow of the corresponding two-dimensional CFT. Our formalism has several benefits compared with other existing frameworks. One can construct the theory or partition function only by using the data of existing $Z_{N}$ symmetric modular invariants (see [8, 9, 10] for example). They usually contain the information of central charge, conformal dimensions, modular $S$ matrix and one can see the type or $Z_{N}$ anomaly of anyons. For example, in the discussion for massless flow, our work provides general data of $Z_{N}$ anomaly classification of (chiral or nonchiral) anyons and this can be thought of as an extension of the pioneering work by Kong and Zheng [21] and older work by Schellekens and Gato-Rivera[38, 39, 40, 41, 42, 43] (We list the reviews by Fuchs and his collaborators for useful references[136, 137]). In the older literature by Schellekens and Gato-Rivera, there seem to exist remarks corresponding to our method (or the situation considered by Kong and Zheng). However, the resultant partition functions are difficult to read. Moreover, as far as we have checked, the general case that we study has not been mentioned in the related literature[136, 137]. Related arguments on nonchiral anyons in some particular TOs can be seen only in [138, 139, 140]  [^2] We also note the simplest cases for the $SU(2)$ models have been studied by Tokiro Numasawa, in his unpublished notes. We thank Tokiro Numasawa for sharing his notes with us.. For the readers interested in the implications of this work in numerical and experimental settings in condensed matter, we remark on four phenomena or models, multilayer or multicomponent system[142], nonchiral anyon in the antiPfaffian and higher Landau level (HLL), and controversy of thermal Hall conductance[134, 135], and Landau level mixing. They can be explained by using our formalism without inconsistencies at this stage. Respectively, quantum field theoretic explanations of such phenomena have already been proposed (see review [143] and reference therein), but we believe our formalism is the most concise, based on the anomaly analysis and anyon physics (or a modern version of quark-hadron particle physics). Moreover, in the discussion of massive RG flow, we provide a new class of BCFT, $Z_{N}$ extended BCFT which determines the gapped phase of $1+1$ dimensional quantum lattice models and the bulk properties of the corresponding $2+1$ dimensional TOs. For example, we demonstrate a concise explanation of $1+1$ dimensional $Z_{2}$ symmetry-protected topological (SPT) phase. Hence, our study provides useful data on both bulk and boundary classifications of TOs.
@@
Compared with the existing literature requiring knowledge of category theory, the requirement of our formalism is much simpler, the table of conformal dimension in a modular invariant. Hence one can obtain the benefit of established systematic construction of coset CFTs[144, 145, 146]. We itemize the fundamental properties to check the data of CFTs or the corresponding modular forms,
@@
The first point determines the fundamental structure of generalized electronic objects, such as Majorana, $Z_{N}$ parafermion. The second point characterizes the stability of anyons, and the third point combined with the fusion algebra of anyons provides the asymptotic behavior of the wavefunctions[51, 54, 74]. Combining methods in this work and the existing matrix-product-state (MPS) or tensor-network calculations [147, 148, 149, 150, 151, 152, 153] and related earlier studies in mathematical physics[18], more general construction of TOs will become possible in principle.
The rest of the manuscript is organized as follows. In section II, we introduce the basics of the bosonic CFTs and their defects and symmetries from a modern perspective. We also provide the fusion algebra data of Ising and Majorana CFTs for later discussions. In section III, we discuss applications of symmetry-based arguments on the classification of massive RG flows from CFTs. We study the implication of noninvertible symmetries and the basic data of the new BCFT corresponding to both the spontaneous symmetry-breaking phase (SSB) and disordered phase. In section IV, we introduce the improved expression of the anomaly-freeness in a CFT and the corresponding TQFT with emphasis on the correspondence between the domain wall and massless RG flow. A new series of modular invariant or domain walls involving nontrivial mixing of chiralities is shown. One can obtain sufficient data from a single modular partition function Eq. (65). Section V is the concluding remarks. We summarize the implications of our formalisms and note general phenomenologies behind gauging procedures which will be useful in establishing a concrete understanding of related phenomena in physics. In the Appendix, we note several detailed systematic constructions of the models in the main texts with historical remarks.

### v3

@@ Abstract
We formulate the $Z_{N}$ group gauging operations (or group extensions) in (smeared) boundary conformal field theories (BCFTs) and bulk conformal field theories, and study their applications to various phenomena in topologically ordered systems. We apply the resultant theories to analyze the correspondence between the renormalization group (RG) flow of CFTs and the classification of topological quantum field theories with the testable information of general classes of partition functions. One can obtain the bulk topological properties of $2+1$ dimensional topological ordered phase corresponding to the massive RG flow of $1+1$ dimensional systems equivalent to the smeared BCFT. We present an obstruction to mass condensation for the smeared BCFT analogous to the Lieb-Shultz-Mattis theorem for noninvertible symmetry. Corresponding to the bulk topological degeneracies in $2+1$ dimensions and gapped quantum phases in $1+1$ dimensions, we construct a new series of BCFT.
We also study the implications of the massless RG flow of $1+1$ dimensional CFTs corresponding to the domain wall or coupling problem in $2+1$ dimensional topological orders which corresponds to the earlier proposal by L. Kong and H. Zheng in [Nucl. Phys. B 966 (2021), 115384], arXiv:1912.01760 closely related to the integer-spin simple current by Schellekens and Gato-Rivera. We reveal that two distinct kinds of massless RG flow exist, coming from anomaly cancellation and anomaly matching related to Witt equivalence of bulk CFTs. The anomaly cancellation flows produce a new series of modular invariants as coupled models, and the anomaly matching flows produce nontrivial domain walls changing the chirality of anyons. Our paper gives a unified direction for the future theoretical and numerical studies of the topological phase based on the established data of classifications of conformal field theories or modular invariants (and the corresponding category theories).
@@
Anyon condensation and Witt equivalence are the most fundamental structures classifying topological orders (TOs) or corresponding topological quantum field theories (TQFTs) in contemporary physics and related mathematics[1, 2, 3]. The theories are deformed to each other without obstruction when the two theories are properly related by the classifications. Whereas the systematic studies on such classification schemes are still in development and sometimes contain controversies, the fundamental significance of abstract algebra of anyons and the corresponding category theories have already been established to some extent (see [4, 5, 6, 7] and reference therein, for example). Combining these ideas with gauging (or group extension) and orbifolding which has been studied in the context of renormalization group (RG) flow[8, 9, 10], one can expect some general correspondence between the classification of TOs in $2+1$ dimensions and that of the RG flows in $1+1$ dimensional CFTs[11, 12, 13, 14, 15]. This is a modern version of bulk-edge correspondence or correspondence between conformal field theory (CFT) and TQFT, known as CFT/TQFT[16, 17].
@@
The symmetry algebraic or category-theoretic version of CFT/TQFT[18, 19, 20, 21, 22] [^1] We thank Liang Kong for notifying us a lot references and their status in the fields., called “topological holography" [24], has been studied widely recently (However, the terminology “topological holography" appeared earier in a different contenxt in [25] representing systems satisyfying both topological and holographic property). The resultant sandwich construction[26, 27, 28] has been applied for several models[29, 30, 31, 32, 33, 34, 35, 36]. Based on these backgrounds, constructing a classification of CFTs is fundamentally important for the classification of TOs. However, we stress that in such generalized symmetry-based studies, the theoretical and mathematical framework on the famous TOs, fractional quantum Hall effects (FQHEs), have still been under development regardless of their experimental significance[37]. It might be surprising for readers, but the existing famous framework, modular-tensor-category (MTC) or bosonic chiral CFT (CCFT), does not correspond to the FQHEs in general. It is necessary to extend the bosonic theory by a finite group and this is called simple current extension[38, 39, 40, 41, 42, 43]. When concentrating our attention on the $Z_{2}$ or fermionic case, such extended structure has already appeared in the established works [17, 44, 45, 46, 47, 48]. More generally, one can see the corresponding arguments widely in older theoretical literature [45, 49, 50, 51, 52, 53, 54]. The resultant structure has captured attention of the fields as realizations of supersymmetry in condensed matter[55, 56, 57, 58, 59, 60, 61, 62, 63]. More recently, the nontriviality of the corresponding categorical frameworks has been studied in [64, 65, 66]. The main difficulty in studying group extensions of CFTs is the treatment of (fermionic) zero modes [67] (or semion[68]). This zero mode counting results in the degeneracies of wavefunctions of FQHEs[17, 69], but the semion structure has not been studied extensively except for several earlier works[67, 46, 44, 70, 71, 72] and the recent works by the author [61, 73, 12, 74, 32]. This is in parallel to the recent development of the studies on fermionic or gauged models[75, 76, 77, 78, 79, 80, 81, 82] (The references are huge, so we have only listed established literature or literature containing discussions on indicators). The zero modes are closely related to the order-disorder duality of lattice models[83], and the group extension is a procedure to unify order and disorder operators canonically in QFTs. Hence, by including nonlocal objects, the theory after the extension becomes outside of MTCs. The resultant theory has been studied as a super-fusion category or premodular fusion category, for example, but this research direction is still in development.
@@
Figure 1: Summary of the relationship between RG flow in CFTs and the notions in TQFTs. The upper figure shows $D$-dimensional TQFT with two boundaries. A FQH system in cylinder geometry is a typical example. The massive RG flow of CFTs or the (extended) smeared BCFTs[84] in this work determine the structure of bulk topological order with a close connection to the entanglement entropy[85, 86, 87, 88, 89]. In TQFT, the massive RG flow in CFTs corresponds to the gluing operation of the two boundaries. On the other hand, the massless RG flow of CFTs implies the (Witt) equivalence relation between group extended TQFTs (or extended SFCs[90, 91]). Applying the interface or domain wall known as a gapped or symmetry-preserving domain wall[11] in TQFT corresponds to the massless RG flow. By analyzing $BCFT_{D}$ with $D\geq 3$ as an ancillary CFT [92], a similar phenomenology will be true in general. We also note that the existing Witt equivalence for MTC (not SFC) seems too strong when applying to realistic settings[93, 6, 94, 95] (we stress that this does not mean their arguments are wrong. This implies the significance of extension for application to more realistic settings). One needs to introduce (or relax) the equivalence relation to the gauged bulk or chiral CFT.
@@
When studying classifications of CFTs, there exist two fundamental aspects: symmetry and renormalization group (RG) flow. Recently, it has been gradually established that a RG flow to a gapped phase can be mapped to the classification of boundary conformal field theory (BCFT)[96, 97, 98, 84], and a massless RG flow [99, 100, 101] can be mapped to the classification of the defect (or interface) of CFTs[102, 103, 104] (We also note the literature on the massless RG flows [105, 106, 107, 108, 109, 110, 111, 112, 113]). In the context of CFT/TQFT, the former corresponds to the classification of bulk states or entanglement spectrum of TOs [85, 86, 87, 88, 89] and the latter corresponds to the classification of boundary anyons induced by the Witt equivalence [11, 12, 13, 14] (Fig. 1). Corresponding discussions for the TQFT side can be seen in the studies of the gapped domain wall in [114, 115, 93, 19, 6, 20, 116, 21, 117] which can be interpreted as a particular class of symmetry-preserving domain walls in [11, 12].
@@
The idea of RG domain wall is surprisingly concise because this is just attaching two systems connected under a massless RG flow [99, 100, 101] through a particular type of domain wall[102, 104]. Hence, one can map the problem of RG to an algebraic problem of a domain wall (or conformal interface) between ultraviolet (UV) and infrared (IR) CFTs. Usually, the verification of RG requires extensive analytical calculations of quantum field theories (QFTs) and numerical tests in lattice models, and they tend to be difficult to formulate rigorously or mathematically. Moreover, by using a folding trick [118] to the system divided by an RG domain wall, one can map the problem to the product of RG-connected CFT with a boundary[102] and this is in the scope of BCFT analysis. Hence, one can apply techniques in defect or boundary CFTs rigorously and systematically in principle (Fig.2). Moreover, the validity of the RG domain wall in the minimal CFTs has been tested more rigorously in [119], and one can expect that the RG domain wall can provide a rigorous formulation of RGs. In this sense, further systematic studies on CFT with defects and boundaries are fundamentally important for the systematic classifications of TOs. We note recent works for the RG domain wall and related junction problems[120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 15, 134].
@@
Figure 2: Folding trick applied to coupled models. By applying the folding trick[118], one can obtain the correspondence between a coupled model ($CFT_{D}\otimes CFT^{\prime}_{D}$) and anyon transformation law of two theories($D_{RG}:CFT_{D}\rightarrow\overline{CFT^{\prime}_{D}}$). Similar arguments can be seen in [11, 117], for example. Because of the folding, the orientation or chirality of one of the two bulk CFTs is reversed. The coupled models in this work are the multicomponent or multilayered TOs. The mapping $D_{\text{RG}}$ corresponding to the symmetry-preserving domain wall[11] or RG domain wall[103, 104] classifies the TOs, possibly including FQHEs with nonchiral anyon. By shrinking one of the two systems, $CFT^{\prime}_{D}$ in the righthand side of the figure, one can obtain the mismatch of bulk theory $CFT_{D}$ and boundary theory $CFT^{\prime}_{D}$. This explains recent controversies on the thermal Hall conductance in experimental settings[135, 136].
@@
In this work, we systematically study algebraic properties of $2+1$ dimensional TOs as the RG flow of the corresponding two-dimensional CFT. Our formalism has several benefits compared with other existing frameworks. One can construct the theory or partition function only by using the data of existing $Z_{N}$ symmetric modular invariants (see [8, 9, 10] for example). They usually contain the information of central charge, conformal dimensions, and modular $S$ matrix, and one can see the type or $Z_{N}$ anomaly of anyons. For example, in the discussion for massless flow, our work provides general data of $Z_{N}$ anomaly classification of (chiral or nonchiral) anyons and this can be thought of as an extension of the pioneering work by Kong and Zheng [21] and older work by Schellekens and Gato-Rivera[38, 39, 40, 41, 42, 43] (We list the reviews by Fuchs and his collaborators for useful references[137, 138]). In the older literature by Schellekens and Gato-Rivera, there seem to exist remarks corresponding to our method (or the situation considered by Kong and Zheng). However, the resultant partition functions are difficult to read. Moreover, as far as we have checked, the general case that we study has not been mentioned in the related literature[137, 138]. Related arguments on nonchiral anyons in some particular TOs can be seen only in [139, 140, 141]  [^2] We also note the simplest cases for the $SU(2)$ models have been studied by Tokiro Numasawa, in his unpublished notes. We thank Tokiro Numasawa for sharing his notes with us.. For the readers interested in the implications of this work in numerical and experimental settings in condensed matter, we remark on four phenomena or models, multilayer or multicomponent system[143], nonchiral anyon in the antiPfaffian and higher Landau level (HLL), and controversy of thermal Hall conductance[135, 136], and Landau level mixing. They can be explained by using our formalism without inconsistencies at this stage. Respectively, quantum field theoretic explanations of such phenomena have already been proposed (see review [144] and reference therein), but we believe our formalism is the most concise, based on the anomaly analysis and anyon physics (or a modern version of quark-hadron particle physics). Moreover, in the discussion of massive RG flow, we provide a new class of BCFT, $Z_{N}$ extended BCFT which determines the gapped phase of $1+1$ dimensional quantum lattice models and the bulk properties of the corresponding $2+1$ dimensional TOs. For example, we demonstrate a concise explanation of $1+1$ dimensional $Z_{2}$ symmetry-protected topological (SPT) phase. Hence, our study provides useful data on both bulk and boundary classifications of TOs.
@@
Compared with the existing literature requiring knowledge of category theory, the requirement of our formalism is much simpler, the table of conformal dimension in a modular invariant. Hence, one can obtain the benefit of the established systematic construction of coset CFTs[145, 146, 147]. We itemize the fundamental properties to check the data of CFTs or the corresponding modular forms,
@@
The first point determines the fundamental structure of generalized electronic objects, such as Majorana, $Z_{N}$ parafermion. The second point characterizes the stability of anyons, and the third point combined with the fusion algebra of anyons provides the asymptotic behavior of the wavefunctions[51, 54, 74]. Combining methods in this work and the existing matrix-product-state (MPS) or tensor-network calculations [148, 149, 150, 151, 152, 153, 154] and related earlier studies in mathematical physics[18], a more general construction of TOs will become possible in principle.
The rest of the manuscript is organized as follows. In section II, we introduce the basics of the bosonic CFTs and their defects and symmetries from a modern perspective. We also provide the fusion algebra data of Ising and Majorana CFTs for later discussions. In section III, we discuss applications of symmetry-based arguments on the classification of massive RG flows from CFTs. We study the implications of noninvertible symmetries and the basic data of the new BCFT corresponding to both the spontaneous symmetry-breaking phase (SSB) and disordered phase. In section IV, we introduce the improved expression of the anomaly-freeness in a CFT and the corresponding TQFT with emphasis on the correspondence between the domain wall and massless RG flow. A new series of modular invariant or domain walls involving nontrivial mixing of chiralities is shown. One can obtain sufficient data from a single modular partition function Eq. (65). Section V is the concluding remarks. We summarize the implications of our formalisms and note general phenomenologies behind gauging procedures which will be useful in establishing a concrete understanding of related phenomena in physics. In the Appendix, we note several detailed systematic constructions of the models in the main texts with historical remarks.
@@
I.1 Summary of the main results
In this manuscript, we provide the construction and analysis of BCFTs, coupled or multicomponent CFTs corresponding to a series of domain walls. Corresponding to these topics, we summarize the main (or new) results in this work.
In section II and III, we provide a general analysis of smeared BCFTs and their implications for $1+1$ dimensional gapped phase. A general condition forbidding the appearance of smeared Cardy states as gapped ground states is discussed. The most essential condition is summarized as Eq. (34) relating the existence of noninvertible symmetry of CFT and condensation conditions in the resultant gapped phase. We also provide a new class of boundary states, eq.(55) and eq.(56), which correspond to an SSB or smeared BCFT.
In section IV, we provide a new series of multicomponent CFTs and corresponding domain walls. The most fundamental finding in this section is summarized in the form of a partition function, Eq. (65). The phenomenologies are summarized in two figures, Fig. 7 and Fig. 9, respectively. For the readers interested in the realization of the multicomponent model, we demonstrate a construction in $1+1$ dimensional quantum lattice models in Appendix B and interpretations of the corresponding wavefunctions in Appendix C. One can obtain a more intuitive understanding of the nontriviality of the mixing of chirality from Fig. 11.